KingPizza:

Last week 35% of the candy sold at a candy store was chocolate. Which model could be used in a simulation conducted to find out whether or not a random customer will choose chocolate? A: a spinner with sectors of 126o and 234o B: tossing a coin 35 times C: a bag of 35 marbles D: a spinner with sectors of 35o and 65o

QuestionCoveBot:

Can you screenshot the models?

snowflake0531:

Okay so we have 35%, which means that the other half is 65% We're doing simulations/models, and 35% of of the 100%, so we can eliminatethe the second and third choice

snowflake0531:

And then for the next step, we see the degrees 35% and 65% are out of 100, which means that we have to convert that percentage to a circle, from 35 out of 100 to blank out of 360 So, this clearly eliminates the last choice too

snowflake0531:

And to check that it is126 and 234, you can do \[\frac{ 35 }{ 100 } = \frac{ x }{ 360 }\] and \[\frac{ 65 }{ 100 } = \frac{ y }{ 360 }\] By this, you'll find the measure 126 and 234

snowflake0531:

So, anyways, you should know which one is the answer now